A composite wire of a uniform cross section 5.5* 10 power minus 5 m square .consists of a steel wire of length 1.5 m and a copper wire of length 2m. by how much will it be stretched when loaded with a mass of 200kg. Young modulus of Steel is 2*10^11 Nm^-2 and copper is 10^11Nm^2.

Grade:11

1 Answers

vanshith
5 years ago
A=5.5*10-5  l=1.5m
PS+pc--->Delta p
Ys=F/A÷deltap/l
deltaps=F.P/A.Ys
           =2000*1.5/5.5*10-5*2*1011
           =3*10-2/55*2
           =0.00027m
Deltapc=2000*20/5.5*10-5*1011
                 =0.00072m
Deltap=PS+pc=1mm